# Hierarchical Uncertainty-Aware Graph Neural Network

### v1

@@ Abstract.
Recent research on graph neural networks (GNNs) has explored mechanisms for capturing local uncertainty and exploiting graph hierarchies to mitigate data sparsity and leverage structural properties. However, the synergistic integration of these two approaches remains underexplored. In this work, we introduce a novel architecture, the Hierarchical Uncertainty-Aware Graph Neural Network (HU-GNN), which unifies multi-scale representation learning, principled uncertainty estimation, and self-supervised embedding diversity within a single end-to-end framework. Specifically, HU-GNN adaptively forms node clusters and estimates uncertainty at multiple structural scales from individual nodes to higher levels. These uncertainty estimates guide a robust message-passing mechanism and attention weighting, effectively mitigating noise and adversarial perturbations while preserving predictive accuracy on both node- and graph-level tasks. We also offer key theoretical contributions, including a probabilistic formulation, rigorous uncertainty-calibration guarantees, and formal robustness bounds. Finally, by incorporating recent advances in graph contrastive learning, HU-GNN maintains diverse, structurally faithful embeddings. Extensive experiments on standard benchmarks demonstrate that our model achieves state-of-the-art robustness and interpretability.
Graph neural networks, adaptive message-passing, local smoothing
@@ 1. Introduction
To solve this limitation, Graph contrastive learning (GCL) frameworks such as Deep Graph Infomax (Velickovic et al., 2019) and Scattering GCL (He and Others, 2024), promote representation diversity by aligning augmented views. However, these methods remain vulnerable to adversarial structural or feature noise. Existing defenses range from preprocessing (Wu et al., 2019) to robust aggregation schemes (Zügner and Günnemann, 2019; Zhang and Others, 2023), yet they typically ignore multi-scale context and offer limited guarantees against adaptive attacks (Xi and Others, 2024; Zhang et al., 2020a). To mitigate these shortcomings, recent work introduces uncertainty-gated message-passing, which retains higher-order expressiveness while attenuating unreliable signals by modulating each message with a learned uncertainty score.
@@
As another branch, recent work has begun to inject uncertainty into GNN pipelines. For example, UAG (Feng et al., 2021) combines Bayesian uncertainty estimation with an attention mechanism, and UnGSL (Han et al., 2025) assigns each node a single confidence score to down-weight edges from low-confidence neighbors, mitigating errors from unreliable signals. However, these flat approaches operate at a single graph resolution and cannot choose when to leverage community-level information, even when immediate neighbors are misleading. Moreover, they do not propagate uncertainty across network layers, leaving deeper reasoning stages unaware of earlier reliability cues. Although local–global hybrids such as LG-GNN (Yu et al., 2024) compute SimRank-style similarities to blend neighborhood and global views, they still ignore uncertainty in predictions. This limitation also affects heterophily-specific models like MixHop (Abu-El-Haija et al., 2019) and FAGCN (Bo et al., 2021), which hard-code higher-order propagation rules without accounting for reliability.
@@
In this paper, we propose a Hierarchical Uncertainty‑aware Graph Neural Network (HU‑GNN) that combines multi‑scale structure with dynamic confidence estimation. HU‑GNN introduces three nested levels of reasoning: (i) a local layer that updates each node by weighting neighboring messages according to feature similarity and learned uncertainty; (ii) a community layer that pools nodes into differentiable clusters, producing super‑node embeddings and an aggregate uncertainty that measures intra‑cluster consensus; and (iii) a global node that summarises the entire graph while tracking global distributional shift. Uncertainty is not a static mask but a latent variable that’s recalculated at every layer and fed back into subsequent propagation, enabling the model to adaptively favor the most reliable scale. In homophilic graphs, HU-GNN effectively becomes a conventional attention-based GNN with enhanced reliability. In heterophilic graphs, it down-weights misleading one-hop edges and instead ascends the hierarchy to community- or global-level contexts.
@@
Beyond its architecture, we deliver new theoretical insights and practical guarantees. First, we derive PAC-Bayesian generalization bounds (Neyshabur et al., 2017) showing that uncertainty gating effectively reduces the effective node degree and yields tighter guarantees. Next, we prove that the joint update of features and uncertainties constitutes a contraction mapping and therefore converges (Forti and Nistri, 2003), stabilizing the hierarchical feedback loop. Finally, we demonstrate that HU-GNN’s error scales only with the uncertainty of misleading neighbors, formally linking robustness to the learned confidence scores under adversarial heterophily (Shanthamallu et al., 2021).
In summary, our contributions can be summarized as,
@@
  • We build an adaptive hierarchy by jointly learning node clusters (communities) and GNN parameters, enabling multi-scale representations while preserving local features. Cluster assignments are iteratively refined as embeddings evolve.
@@
  • Inspired by energy-based modeling, we estimate uncertainties at multiple scales (local, cluster, and global) with corresponding uncertainty scores. These scores reflect how anomalous or unpredictable each node or cluster is relative to the learned distributions.
@@
  • We leverage these uncertainty scores for robust aggregation by modifying the message-passing scheme to down-weight neighbors or clusters with high uncertainty, thus preventing adversarial or noisy inputs from corrupting the aggregation.
@@
  • Comprehensive evaluations on real-world benchmarks against leading baselines demonstrate substantial performance gains, corroborating the effectiveness of our approach and validating our theoretical analysis.
@@ 3.1. Problem Formulation
We consider a graph $G=(|\mathcal{V}|,|\mathcal{E}|)$ with node set $|\mathcal{V}|=n$ and edge set $|\mathcal{E}|=m$. The structural property of $\mathcal{G}$ is represented by its adjacency matrix $A\in\{0,1\}^{n\times n}$. A diagonal matrix $D$ of node degrees is derived from $A$ as $d_{ii}=\sum^{n}_{j=1}{A_{ij}}$. Each node $i\in|\mathcal{V}|$ has a feature vector ${x}_{i}\in\mathbb{R}^{d}$. In semi-supervised node classification, a subset of nodes $|\mathcal{V}|_{L}\subset|\mathcal{V}|$ has observed labels $y_{i}$ (e.g. research paper topics in a citation network), and we aim to predict labels for the unlabeled nodes $|\mathcal{V}|\setminus|\mathcal{V}|_{L}$.
@@ 3.2. Graph Homophily
Homophilic graphs have a high probability that an edge connects nodes of the same class, while heterophilic graphs often connect nodes of different classes. A homophily ratio $\mathcal{H}$ close to 1 means strong homophily as follows:
(1) $\mathcal{H}\equiv\frac{\sum_{(i,j)\in\mathcal{E}}1(Y_{i}=Y_{j})}{|\mathcal{E}|}$
Traditional GNNs (Kipf and Welling, 2016; Velickovic et al., 2017) perform poorly when $\mathcal{H}$ is low, because aggregating information from a predominantly dissimilar neighborhood confuses the classifier. Recent heterophilic GNN designs address this by extending the neighborhood definition or by refining the aggregation to distinguish between similar and dissimilar neighbors.
@@ 3.3. Message-Passing Scheme
Here, $H^{(0)}=X$ denotes the original feature matrix, and $H^{(l)}$ represents the hidden states at layer $l$. The function $\sigma$ (e.g., ReLU) introduces nonlinearity, while $W^{(l)}$ is a layer-specific weight matrix shared across all nodes. After $L$ layers, the model’s output $H^{(L)}$ is fed into a negative log-likelihood loss to produce predictions:
(3) $\widehat{Y}=\mathcal{L}_{\mathrm{nll}}\bigl{(}H^{(L)}\bigr{)},$
and the parameters are optimized by minimizing $\mathcal{L}_{\mathrm{nll}}\bigl{(}\widehat{Y},Y\bigr{)}$ against the ground-truth labels $Y$. Since many GNN architectures assume a homophilic structure, they inherently smooth node features by emphasizing low-frequency components of the graph signal (Nt and Maehara, 2019).
@@ 3.4. Uncertainty Types
We incorporate two broad types of uncertainty in our model: (1) epistemic (model) uncertainty and (2) aleatoric (data) uncertainty. Epistemic uncertainty reflects uncertainty in the model parameters or structure. In our context, we consider uncertainty about the graph connectivity or the GNN weights due to limited training sets (semi-supervised learning). Aleatoric uncertainty reflects inherent noise in the data. For example, a paper that cites very diverse other papers might have ambiguous features, or a node’s label might be difficult to predict even with full information. HU-GNN uses a Bayesian-inspired approach to handle both: it maintains distributions over node representations and uses the spread (variance) of these distributions as a measure of uncertainty. At the local level, uncertainty might come from a neighbor’s feature noise or an edge that is possibly spurious; at the group level, uncertainty can arise if a community’s internal consensus is low (the members have widely varying features or labels); at the global level, uncertainty could stem from distribution shift or class imbalance in the entire graph (e.g. if some classes are under-represented, predictions across the graph for that class are less certain).
@@ 4. Methodology
    Local Message-Passing: Each node aggregates information from its neighbors, weighting each by its uncertainty or reliability.
@@
    Community Pooling: Assume a higher-order grouping of nodes into communities. Each community is treated as a super-node with its embedding and uncertainty. This approach captures higher-order structures beyond immediate links and offers improved context by providing clues from more distant nodes.
@@
    Global Integration We introduce a global context node connected to every community. It aggregates an overall representation, capturing aspects like class proportions or a feature summary, and maintains a global uncertainty. The global node passes its message to individual nodes.
@@
${x}_{i}\in\mathbb{R}^{d}$ Raw feature vector of node $i$
@@
Figure 1. Overall framework of HU-GNN, illustrating the classification scenario for node $i$. Dotted lines indicate low confidence, while solid lines indicate high confidence
@@ 4.1. Local Message-Passing
This is the bottom (node-level) layer which updates each node’s embedding by aggregating messages from its one-hop neighbors, using learned weights that account for neighbor uncertainties. We initialize each node’s embedding ${h}_{i}^{(0)}={x}_{i}$ (the raw features or an MLP-transformed feature) and an initial uncertainty $u_{i}^{(0)}$ for each node (more on initialization later). For notational brevity, let us assume the projection of node features as below:
(4) $\tilde{h}^{(1)}_{i}={W}_{O}^{(1)}{h}_{i}^{(0)},$
where ${W}_{O}^{(1)}$ is a trainable weight matrix for local message-passing. The local aggregation can be defined as:
(5) ${h}_{i}^{(1)}=\sigma\Big{(}\tilde{h}^{(1)}_{i}+\sum_{j\in N(i)}m_{ij}^{(1)}\tilde{h}^{(1)}_{j}\Big{)},$
and $\sigma(\cdot)$ is an activation function (e.g. ReLU). The crucial part is the neighbor weight $m_{ij}^{(1)}$, which we make feature- and uncertainty-aware. One simple choice is to let $m_{ij}^{(1)}$ consider both feature similarity and uncertainty as below:
(6) $m_{ij}^{(1)}=\frac{1}{2}\left(\frac{\exp(a_{ij}^{(1)})}{\sum_{k}\exp(a_{ik}^{(1)})}+\frac{\exp(-u_{j}^{(1)})}{\sum_{k}\exp(-u_{k}^{(1)})}\right),\;\;k\in N(i)$
Here, $\alpha^{(1)}_{ij}=\vec{a}^{T}\left(\tilde{h}^{(1)}_{i}\;\|\;\tilde{h}^{(1)}_{j}\right)$ is an attention value (please refer to GAT (Velickovic et al., 2017) for details) and $u^{(1)}$ represents uncertainty level. Thus, node $j$ with higher prior uncertainty or lower attention score will have a smaller influence on $i$’s update. This improves the UnGSL approach (Han et al., 2025), which reflects both feature similarity and uncertainty. In summary, Eq. 6 produces a new embedding ${h}_{i}^{(1)}$ as a feature- and uncertainty-weighted average of neighbor information.
Next, we introduce how to update the node’s uncertainty $u_{i}$ after this aggregation. Intuitively, if node $i$ received consistent messages from multiple confident neighbors, its uncertainty should decrease; if the messages were conflicting or from uncertain neighbors, its uncertainty should remain high or even increase. We formalize one way to update uncertainty as:
(7) $u_{i}^{(1)}=f_{u}\left(\frac{1}{|N(i)|}\sum_{j\in N(i)}\|\tilde{h}^{(1)}_{i}-\tilde{h}^{(1)}_{j}\|^{2}\right),$
Here $f_{u}(\cdot)$ is a function that computes the new uncertainty of node $i$ based on its prior uncertainty and the uncertainties and features of its neighbors. A possible choice for $f_{u}$ is to use the variance of the incoming messages (Sihag et al., 2022). If neighbor messages disagree (high variance), $u_{i}^{(1)}$ will be high; if they are very consistent and $i$ was not too uncertain to start, the uncertainty will be lower. In practice, using the Sigmoid function, one could implement $f_{u}(\cdot)$ as a small neural network taking node features and uncertainties as input and generating an output uncertainty ($0\leq u_{i}\leq 1$). This completes the local layer: each node now has an updated embedding ${h}_{i}$ and uncertainty $u_{i}$, where we set $L=2$ in this paper.
4.2. Community Pooling

### v2

@@ Abstract.
Recent research on graph neural networks (GNNs) has explored mechanisms for capturing local uncertainty and exploiting graph hierarchies to mitigate data sparsity and leverage structural properties. However, the synergistic integration of these two approaches remains underexplored. This work introduces a novel architecture, the Hierarchical Uncertainty-Aware Graph Neural Network (HU-GNN), which unifies multi-scale representation learning, principled uncertainty estimation, and self-supervised embedding diversity within a single end-to-end framework. Specifically, HU-GNN adaptively forms node clusters and estimates uncertainty at multiple structural scales from individual nodes to higher levels. These uncertainty estimates guide a robust message-passing mechanism and attention weighting, effectively mitigating noise and adversarial perturbations while preserving predictive accuracy on semi-supervised classification tasks. We also offer key theoretical contributions, including a probabilistic formulation, rigorous uncertainty-calibration guarantees, and formal robustness bounds. Extensive experiments on standard benchmarks demonstrate that our model achieves state-of-the-art robustness and interpretability.
Graph neural networks, graph heterophily, hierarchical structure learning, uncertainty-aware message-passing, PAC-Bayes generalization bounds
@@ 1. Introduction
As one solution, graph contrastive learning frameworks such as Deep Graph Infomax (Velickovic et al., 2019) and Scattering GCL (He and Others, 2024), promote representation diversity by aligning augmented views. However, these methods remain vulnerable to adversarial structural or feature noise. Existing defenses range from preprocessing (Wu et al., 2019) to robust aggregation schemes (Zügner and Günnemann, 2019; Zhang and Others, 2023), yet they typically ignore multi-scale context and offer limited guarantees against adaptive attacks (Xi and Others, 2024; Zhang et al., 2020a). Recent work introduces uncertainty-gated message-passing to mitigate these shortcomings, which retains higher-order expressiveness while attenuating unreliable signals by modulating each message with a learned uncertainty score.
@@
As another branch, recent work has begun to inject uncertainty into GNN pipelines. For example, UAG (Feng et al., 2021) combines Bayesian uncertainty estimation with an attention mechanism, and UnGSL (Han et al., 2025) assigns each node a single confidence score to down-weight edges from low-confidence neighbors, mitigating errors from unreliable signals. However, these flat approaches operate at a single graph resolution and cannot decide when to leverage community‑level information, even when immediate neighbors are misleading. Moreover, they do not propagate uncertainty across network layers, leaving deeper reasoning stages unaware of earlier reliability cues. Although local–global hybrids such as LG-GNN (Yu et al., 2024) compute SimRank-style similarities to blend neighborhood and global views, they still ignore uncertainty in predictions. This limitation also affects heterophily-specific models like MixHop (Abu-El-Haija et al., 2019) and FAGCN (Bo et al., 2021), which hard-code higher-order propagation rules without accounting for reliability (or low confidence).
@@
In this paper, we propose a Hierarchical Uncertainty‑aware Graph Neural Network (HU‑GNN) that combines multi‑scale structure with dynamic confidence estimation. Specifically, HU‑GNN introduces three nested levels of reasoning: (i) a local layer that updates each node by weighting neighboring messages according to feature similarity and learned uncertainty; (ii) a community layer that pools nodes into differentiable clusters, producing super‑node embeddings and an aggregate uncertainty that measures intra‑cluster consensus; and (iii) a global node that summarises the entire graph while tracking global distributional shift. Uncertainty is not a static mask but a latent variable that’s recalculated at every layer and fed back into subsequent propagation, enabling the model to adaptively favor the most reliable scale. In homophilic graphs, HU-GNN effectively becomes a conventional attention-based GNN with enhanced reliability. In heterophilic graphs, it down-weights misleading one-hop neighbors and instead ascends the hierarchy to community- or global-level contexts.
@@
Beyond its architecture, we deliver new theoretical insights and practical guarantees. First, we derive PAC-Bayesian generalization bounds (Neyshabur et al., 2017) showing that uncertainty gating effectively reduces the effective node degree and yields tighter guarantees. Next, we prove that the joint update of features and uncertainties constitutes a contraction mapping and converges (Forti and Nistri, 2003), thereby stabilizing the hierarchical feedback loop. Finally, we demonstrate that the error of HU-GNN scales only with the uncertainty of misleading neighbors, formally linking robustness to the learned confidence scores under adversarial heterophily (Shanthamallu et al., 2021).
In summary, our contributions can be summarized as follows:
@@
  • We construct an adaptive hierarchy by jointly learning node clusters and GNN parameters, enabling multi-scale representations while preserving local feature information. Cluster assignments are refined iteratively as the embeddings evolve.
@@
  • Inspired by energy-based modeling, we estimate uncertainties at multiple scales (local, community, and global) using corresponding uncertainty scores. These scores quantify how anomalous or unpredictable each node or cluster is relative to the learned distributions.
@@
  • We utilize these uncertainty scores to achieve robust aggregation by modifying the message-passing scheme, down-weighting neighbors or clusters with high uncertainty and low attention. This prevents adversarial or noisy inputs from contaminating the aggregation process.
@@
  • Comprehensive evaluations on real-world benchmarks against state-of-the-art baselines show substantial performance improvements, corroborating the effectiveness of our proposed approach and validating our theoretical analysis.
@@ 3.1. Problem Formulation
We consider a graph $G=(|\mathcal{V}|,|\mathcal{E}|)$ with node set $|\mathcal{V}|=n$ and edge set $|\mathcal{E}|=m$. The structural property of $\mathcal{G}$ is represented by its adjacency matrix $A\in\{0,1\}^{n\times n}$. A diagonal matrix $D$ of node degrees is derived from $A$ as $d_{ii}=\sum^{n}_{j=1}{A_{ij}}$. Each node $i\in|\mathcal{V}|$ has a feature vector ${x}_{i}$. In semi-supervised node classification, a subset of nodes $|\mathcal{V}|_{L}\subset|\mathcal{V}|$ has observed labels $y_{i}$ (e.g. research paper topics in a citation network), and we aim to predict labels for the unlabeled nodes $|\mathcal{V}|\setminus|\mathcal{V}|_{L}$.
@@ 3.2. Graph Homophily
Homophilic graphs exhibit a high probability that an edge connects nodes belonging to the same class, whereas heterophilic graphs frequently link nodes from different classes. A homophily ratio $\mathcal{H}$ close to 1 indicates strong homophily and is defined as follows:
(1) $\mathcal{H}\equiv\frac{\sum_{(i,j)\in\mathcal{E}}1(Y_{i}=Y_{j})}{|\mathcal{E}|}.$
Many existing GNN architectures implicitly assume a homophilic structure, where they inherently smooth node features by emphasizing low-frequency components of the graph signal (Nt and Maehara, 2019). Consequently, traditional GNNs (Kipf and Welling, 2016; Velickovic et al., 2017) exhibit suboptimal performance when the homophily ratio $\mathcal{H}$ is low, as aggregating information from predominantly dissimilar neighborhoods can mislead the classifier. Recent approaches for heterophilic GNNs address this limitation either by broadening the neighborhood definition or by refining aggregation schemes to distinguish effectively between similar and dissimilar neighbors during message-passing.
@@ 3.3. Message-Passing Scheme
Here, $H^{(0)}=X$ denotes the original feature matrix, and $H^{(l)}$ represents the hidden states at layer $l$. The function $\sigma$ (e.g., ReLU) introduces nonlinearity, while $W^{(l)}$ is a layer-specific weight matrix shared across all nodes. After $L$ layers, the softmax function $\phi$ is applied to the model’s output $H^{(L)}$ for prediction as below:
(3) $\widehat{Y}=\phi\bigl{(}H^{(L)}\bigr{)},$
and the parameters are optimized by minimizing the negative log-likelihood $\mathcal{L}_{\mathrm{nll}}\bigl{(}\widehat{Y},Y\bigr{)}$ against the ground-truth labels $Y$.
@@ 3.4. Uncertainty Types
We incorporate two broad types of uncertainty in our work: (1) epistemic (model) uncertainty and (2) aleatoric (data) uncertainty. Epistemic uncertainty reflects uncertainty in the model parameters or structure. In our context, we consider uncertainty about the graph connectivity or the GNN weights due to limited training sets (semi-supervised learning). Aleatoric uncertainty reflects inherent noise in the data. For example, a paper that cites very diverse topics might have ambiguous features, or a node’s label might be difficult to predict even with full information. Our method (HU-GNN) uses a Bayesian-inspired approach to handle both: it maintains distributions over node representations and uses the spread (variance) of these distributions as a measure of uncertainty. At the local level, uncertainty might come from a neighbor’s feature noise or an edge that is possibly spurious; at the group level, uncertainty can arise if a community’s internal consensus is low (the members have widely varying features or labels); at the global level, uncertainty could stem from distribution shift or class imbalance in the entire graph (e.g. if some classes are under-represented, predictions across the graph for that class are less certain).
@@ 4. Methodology
    Local Message-Passing: Each node aggregates information from its neighbors, weighting each by its uncertainty and feature-based similarity.
@@
    Community Assignment and Pooling: Assume a higher-order grouping of nodes into communities. Each community is treated as a super-node with its embedding and uncertainty. This approach captures higher-order structures beyond immediate links and offers improved context by providing clues from more distant nodes.
@@
    Global Integration We introduce a global context node connected to every community. It aggregates an overall representation, capturing aspects like class proportions or a feature summary, and maintains a global uncertainty. The global node is used for classification purposes at the end.
@@
${x}_{i}$ Raw feature vector of node $i$
@@
Figure 1. The overall framework of HU-GNN, illustrating message‑passing and pooling for node $i$. Dashed lines indicate connections with low confidence, whereas solid lines represent connections with high confidence
@@ 4.1. Local Message-Passing
The node‑level layer updates each node’s embedding by aggregating messages from its one‑hop neighbors with weights that account for neighbor uncertainties. We initialize each node’s embedding as $h_{i}^{(0)}=x_{i}$ (raw feature vector), and $u_{i}^{(0)}$ is the initial uncertainty (see Section 4.5). We then project the features for brevity as,
(4) $\tilde{h}_{i}^{(1)}=W_{O}^{(1)}\,h_{i}^{(0)},$
where $\tilde{h}_{i}^{(1)}\in\mathbb{R}^{F}$ and $W_{O}^{(1)}\in\mathbb{R}^{F^{\prime}\times F}$ is a trainable weight matrix.
Uncertainty‑aware attention.
We compute neighbor weights by combining feature similarity and node uncertainty:
(5) $m_{ij}^{(1)}=\frac{\exp\bigl{(}a^{\top}[\tilde{h}_{i}^{(1)}\,\|\,\tilde{h}_{j}^{(1)}]\bigr{)}\,\exp\bigl{(}-\,u_{j}^{(1)}\bigr{)}}{\sum_{k\in N(i)}\exp\bigl{(}a^{\top}[\tilde{h}_{i}^{(1)}\,\|\,\tilde{h}_{k}^{(1)}]\bigr{)}\,\exp\bigl{(}-\,u_{k}^{(1)}\bigr{)}},$
where $a\in\mathbb{R}^{2F}$ is the attention vector and $u_{j}^{(1)}\in\mathbb{R}^{1}$ is node $j$’s uncertainty level. Therefore, nodes with higher uncertainty or lower attention receive smaller weights.
Low‑/high‑frequency decomposition.
To handle graph heterophily, we extract low‑frequency (LF) and high‑frequency (HF) components of the aggregated messages (Wu et al., 2023):
(6) $\underbrace{l_{i}^{(1)}}_{\text{LF}}=\sum_{j\in N(i)}m_{ij}^{(1)}\,\tilde{h}_{j}^{(1)},\quad\underbrace{d_{i}^{(1)}}_{\text{HF}}=\tilde{h}_{i}^{(1)}-l_{i}^{(1)}.$
Aggregation.
The node embedding is updated by applying a weighted fusion of LF and HF signals:
(7) $h_{i}^{(1)}=\sigma\Bigl{(}\tilde{h}_{i}^{(1)}+p^{(1)}_{i}\,l_{i}^{(1)}\,+(1-p^{(1)}_{i})\,d_{i}^{(1)}\Bigr{)},$
where $p^{(1)}_{i}=\text{cos}(\tilde{h}_{i}^{(1)},l_{i}^{(1)})$ is a cosine similarity and $\sigma(\cdot)$ is a nonlinear activation (e.g., ReLU). Next, we describe the update rule for each node’s uncertainty $u_{i}$ following the aggregation step. Intuitively, when node $i$ receives consistent messages from multiple low‑uncertainty neighbors, its uncertainty should decrease; conversely, conflicting or uncertain neighbor messages should maintain or increase its uncertainty. One possible formulation is,
(8) $u_{i}^{(1)}=f_{u}\Bigl{(}\frac{1}{|N(i)|}\sum_{j\in N(i)}\bigl{\lVert}\tilde{h}_{i}^{(1)}-\tilde{h}_{j}^{(1)}\bigr{\rVert}^{2}\Bigr{)},$
where $f_{u}(\cdot)$ maps the mean squared deviation of incoming messages to a new uncertainty estimate. For example, $f_{u}$ may compute the sample variance of neighbor features as in (Sihag et al., 2022). Thus, high variance yields a larger $u_{i}^{(1)}$, whereas low variance and a low prior uncertainty lead to a smaller $u_{i}^{(1)}$. In practice, $f_{u}$ can be implemented as a lightweight neural network with a sigmoid activation to enforce $u_{i}^{(1)}\in[0,1]$. This completes the local layer: each node obtains both an updated embedding $h_{i}^{(1)}$ and an updated uncertainty $u_{i}^{(1)}$. In this paper, we set the total number of layers as $l=2$.
Remark. With the slight abuse of notation, we let $h_{i}\,\equiv\,h^{(l=2)}_{i}$ stands for the output of two local message‑passing layers.
4.2. Community Assignment and Pooling